1. The average number of milligrams (mg) of cholesterol in a cup of a certain brand of ice cream is 660 mg, the standard deviation is 35 mg. Assume the variable is normally distributed.


a. If a cup of ice cream is selected, what is the probability that the cholesterol content will be more than 670 mg?



b. If a sample of 10 cups of ice cream is selected, what is the probability that the mean of the sample will be larger than 670 mg?



2.In a study of the life expectancy of 400 people in a certain geographic region, the mean age at death was 70 years, and the standard deviation was 5.1 years. If a sample of 50 people from this region is selected, what is the probability that the mean life expectancy will be less than 68 years?

1
Expert's answer
2022-04-01T04:50:37-0400

1. We have a normal distribution, "\mu=600, \sigma=35."


a.

Let's convert it to the standard normal distribution:

"z=\cfrac{x-\mu}{\sigma}=\cfrac{670-600}{35}=2,"


"P(X>670)=1-P(X<670)=\\
=1-P(Z<2)="

"=1-0.9772=0.0228" ​(from z-table)


b.

Let's convert it to the standard normal distribution,

"z=\cfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\cfrac{670-600}{35/\sqrt{10}}=6.32."

"P(\bar{X}>670)=1-P(\bar{X}<670)=\\
=1-P(Z<6.32)="

"=1-1=0" ​(from z-table)


2. We have a normal distribution, "\mu=70, \sigma=5.1,n=50."

Let's convert it to the standard normal distribution,

"z=\cfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\cfrac{68-70}{5.1/\sqrt{50}}=-2.77."

"P(\bar{X}<68)=P(Z<-2.77)=0.0028" (from z-table)